OpenStudy (ebonyserrano):

Three consecutive integers have a sum of 81. Find the integers. so it would be 26,27 and 81?

OpenStudy (ebonyserrano):

@steve816

OpenStudy (farmboy813):

yes that is correct !!

OpenStudy (crazylizard102):

yes

OpenStudy (ebonyserrano):

thanks!

OpenStudy (ebonyserrano):

it's actually wrong :'(

OpenStudy (ebonyserrano):

I think 81 is wrong

OpenStudy (farmboy813):

omg im so sorry !

OpenStudy (ebonyserrano):

its ok

OpenStudy (ebonyserrano):

can you guys help though?

OpenStudy (thefatdog):

81 = x + x +1 +x + 2 given 81 = 3x + 3 combine like terms -3 -3 subtraction property of equality 78 = 3x /3 /3 division property of equality 26 = x symmetry property x = 26 solution x =26 x + 1 = 27 x+2 = 28\ 26, 27, 28 26 +27 + 28 = 81 so, 26, 27, and 28.